# Pychop: Emulating Low-Precision Arithmetic in Numerical Methods and Neural Networks

### v4

@@ Abstract.
Motivated by the growing demand for low-precision arithmetic in computational science, we exploit lower-precision emulation in Python—widely regarded as the dominant programming language for numerical analysis and machine learning. Low-precision training has revolutionized deep learning by enabling more efficient computation and reduced memory and energy consumption while maintaining model fidelity. To better enable numerical experimentation with and exploration of low precision computation, we developed the Pychop library, which supports customizable floating-point formats and a comprehensive set of rounding modes in Python, allowing users to benefit from fast, low-precision emulation in numerous applications. Pychop also introduces interfaces for both PyTorch and JAX, enabling efficient low-precision emulation on GPUs for neural network training and inference with unparalleled flexibility.
In this paper, we offer a comprehensive exposition of the design, implementation, validation, and practical application of Pychop, establishing it as a foundational tool for advancing efficient mixed-precision algorithms. Furthermore, we present empirical results on low-precision emulation for image classification and object detection using published datasets, illustrating the sensitivity of the use of low precision and offering valuable insights into its impact. Pychop enables in-depth investigations into the effects of numerical precision, facilitates the development of novel hardware accelerators, and integrates seamlessly into existing deep learning workflows. Software and experimental code are publicly available at https://github.com/inEXASCALE/pychop.
@@
ccs: Mathematics of computing Mathematical software performanceccs: Computing methodologies Modeling and simulationccs: Software and its engineering
@@ 1. Introduction
The increasing support of low-precision arithmetic in hardware architectures has triggered a resurgence of interest in mixed-precision algorithms, particularly within the fields of numerical analysis and deep learning. There are numerous benefits to using numerical formats with lower precision than single precision primarily because they require less memory bandwidth. The mixed-precision algorithms, which strategically combine low-precision and high-precision computations, have emerged as a heating topic of research due to their capability to enhance algorithmic performance across various domains—most notably energy efficiency, data transfer, and arithmetic speedup—while maintaining acceptable levels of numerical accuracy and stability (Higham and Mary, 2022). This paradigm exploits the inherent trade-offs between computational cost and precision, offering a compelling technique for addressing the escalating demands of large-scale numerical simulations and machine learning applications. The ability to tailor precision to specific computational tasks not only reduces energy consumption but also reduces processing times, making mixed-precision computations indispensable in the era of exascale computing and resource-constrained environments.
It is known that the increasing size of neural networks typically improves model generalization ability and accuracy at the cost of memory and compute resources for model deployment. The advance of mixed-precision training has been a milestone for deep learning efficiency. Micikevicius et al. (2018) established its use in practice, demonstrating that fp16 (half-precision, 16 bits total with 5 exponent and 10 significand bits) could handle weights and activations while fp32 accumulations preserved gradient fidelity, yielding speedups of 23×2-3$\times$2 - 3 × on NVIDIA Volta GPUs. This approach has been integrated into PyTorch (Paszke et al., 2019) and implemented via Automatic Mixed Precision (AMP) [^1] https://pytorch.org/docs/stable/amp.html, which automates training in low precision with gradient scaling, lowering memory consumption by approximately 50% for models like ResNet-50 (He et al., 2016). However, AMP’s dependence on hardware-supported fp16 hinders its flexibility to other floating-point formats, a constraint our emulator eliminates by supporting arbitrary precision configurations. Further, (Perez et al., 2023) propose the training method via the scalings for fp8 (8-bit floating-point with two formats (Noune et al., 2022; Micikevicius et al., 2022)—exponent bits and 2 significand bits for E5 and 4 exponent bits and 3 significand bits for E4) linear layers by dynamically updating per-tensor scales for the weights, gradients and activations, validating an acceptable performance of GPT and Llama 2 deployed with fp8 precision across model sizes ranging from  [^1] M to 70B.
@@
Table 1. Key parameters of six floating-point formats; $u$ denotes the unit roundoff corresponding to the precision, $x_{\min}$ denotes the smallest positive normalized floating-point number, $x_{\max}$ denotes the largest floating-point number, $t$ denotes the number of binary digits in the significand (including the implicit leading bit), $e_{\min}$ denotes exponent of $x_{\min}$, $e_{\max}$ denotes exponent of $x_{\max}$, with added columns for exponent bits and significand bits (excluding implicit bit).
$\quad u$ $\quad x_{\min}$ $\quad x_{\max}$ $t$ $e_{\min}$ $e_{\max}$ exp. bits sig. bits
NVIDIA quarter precision (q43) $6.25\times 10^{-2}$ $1.5625\times 10^{-2}$ $2.40\times 10^{2}$ 4 -6 7 4 3
NVIDIA quarter precision (q52) $1.25\times 10^{-1}$ $6.10\times 10^{-5}$ $5.73\times 10^{4}$ 3 -14 15 5 2
@@
TensorFloat-32 (tf32) $9.77\times 10^{-4}$ $1.18\times 10^{-38}$ $1.70\times 10^{38}$ 11 -126 127 8 10
@@
We design a low-precision emulation software module called pychop for Python, offering a highly customizable setting for users to simulate arbitrary low-precision arithmetic—floating-point, fixed-point, and integer—with a diverse of rounding modes. This emulator transcends the constraints of fixed hardware by allowing users to define custom precision formats—specifying exponent and significand bits—and to select from rounding modes, e.g., round to nearest, round up / down, round toward zero, round toward odd, and two stochastic variants (proportional and uniform probability). The software promotes comprehensive precision format support, flexibility, and a beginner-friendly interface:
  • Unmatched Flexibility: Emulating both standard (see Table 1) and customized numerical formats for low-precision emulation, enabling researchers to prototype hypothetical hardware or explore theoretical precision boundaries without physical constraints.
  • Precision Granularity: Providing precise control over numerical representation, which is critical for dissecting the impact of quantization on both numerical algorithms and deep learning methods, e.g., gradient updates, weight distributions, and activation ranges in neural networks.
  • Seamless Integration: Offering direct emulators into PyTorch layers, enabling practitioners to experiment with mixed precision in training and inference pipelines with minimal overhead. The user-friendly API enables users to deploy quantization-aware and post-quantization strategies easily.
  • Rounding Mode Exploration: Supporting a diverse collection of rounding modes is available, enabling empirical evaluation of their roles in numerical stability.
@@ 2. Related Work
Table 2. Software for Simulating Low-precision Arithmetic
@@
chop (Higham and Pranesh, 2019) MATLAB fp32/fp64 R S F fp16/bf16 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Pychop Python / MATLAB fp32/fp64 R S F fp16/bf16 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Low-precision arithmetic has emerged as a pivotal technique for optimizing computational efficiency in scientific computing and machine learning applications, where reduced precision can significantly lower resource demands while maintaining acceptable accuracy (Higham and Pranesh, 2019). Several software libraries have been developed to emulate low-precision arithmetic, each with distinct capabilities tailored to specific use cases. In this section, we discuss the software packages listed in Table $2$The table’s design follows (Fasi and Mikaitis, 2023), highlighting their strengths and limitations in the context of low-precision arithmetic simulation.
@@
GNU MPFR (Fousse et al., 2007) is a C library for simulating multiple-precision floating-point computations with guaranteed correct rounding. It excels in scenarios requiring arbitrary precision for both the significand ($p$) and exponent range ($e$). It is a preferred choice for numerical simulations demanding high accuracy, such as symbolic computation and numerical analysis. However, its default lack of subnormal number support ($s$, denoted as O) requires explicit user configuration, which can complicate workflows. Besides, GNU MPFR does not offer built-in floating-point formats and is not suited for neural network training, limiting its capability in low-precision emulation in deep learning.
@@
SIPE (Lefèvre, 2013), another C-based library, focuses on very low-precision computations with correct rounding. It supports multiple storage formats, restricted significand ($p$), and a restricted exponent range ($e$), while fully supporting subnormal numbers ($s$). SIPE implements RNE and RZ rounding modes for low-precision simulations which is suitable for exploring numerical stability and precision trade-offs in low-bitwidth computations, such as those encountered in embedded systems. Its primary limitation lies in its restricted rounding mode support and lack of built-in formats, which reduces its versatility.
@@
rpe (Dawson and Düben, 2017), implemented in Fortran, is tailored for emulating reduced floating-point precision in large-scale numerical simulations, such as climate modeling. It operates with restricted significand ($p$) and exponent range ($e$) limited to built-in types (B) and supports subnormal numbers only for built-in types (B). A key advantage is its native support for the fp16 format, aligning with low-precision hardware standards like IEEE 754 binary16. However, rpe’s exclusive support for RNE rounding limits its flexibility in scenarios requiring diverse rounding strategies, and it does not support neural network training, focusing solely on numerical simulations.
@@
FloatX (Flegar et al., 2019) and FlexFloat (Tagliavini et al., 2020) are C++ libraries that provide frameworks for customized floating-point arithmetic in low-precision simulations. These libraries support restricted significand ((p)) and exponent range ((e)), fully accommodate subnormal numbers ((s)), and utilize fp32/fp64 storage formats for compatibility with standard representations. Their simplicity, limited to round-to-nearest-even (RNE) rounding, makes them accessible for educational purposes and prototyping. However, this restricted rounding support limits their adaptability, and neither library includes predefined formats or supports neural network training, confining their use to general-purpose low-precision arithmetic experimentation. Both FloatX and FlexFloat strictly follow standard C++ conventions, including round-to-nearest, ties-to-even rounding, and default datatype casting. Although these conventions ensure compatibility with native floating-point operations, they limit users’ ability to explore diverse rounding strategies in numerical simulations. This constraint poses a notable challenge for researchers and practitioners needing flexible, application-specific rounding behaviors.
@@
INTLAB (Rump, 1999), a MATLAB toolbox, leverages interval arithmetic to facilitate low-precision floating-point simulation. It uses fp64 storage, supports restricted significand ($p$) and exponent range ($e$), and fully supports subnormal numbers ($s$). INTLAB provides RNE, RZ, and RUD rounding modes, offering moderate flexibility for numerical computations in MATLAB environments, such as verified computing. Its lack of built-in formats and optimization for neural network training limits its scope, positioning it as a tool for reliable numerical analysis rather than machine learning.
@@
chop (Higham and Pranesh, 2019), another MATLAB library, enables low-precision arithmetic simulation with fp32/fp64 storage, restricted significand ($p$), and exponent range ($e$), alongside flexible subnormal number support (F). It supports built-in fp16 and bf16 formats, aligning with modern hardware standards, and implements RNE, RZ, RUD, and stochastic rounding (SR). The inclusion of SR is particularly valuable for simulating quantization effects, which are critical in machine learning research. Despite this, chop does not directly support neural network training and inference, and it is limited to the MATLAB environment, restricting its application to numerical experimentation and analysis of quantization impacts in machine learning.
@@
CPFloat (Fasi and Mikaitis, 2023) is a C library optimized for efficient low-precision arithmetic simulation, supporting fp32/fp64 storage with restricted significand ($p$), and exponent range ($e$). It includes built-in fp16, bf16, and tf32 formats with flexible subnormal number support (F), enhancing compatibility with hardware-accelerated systems such as GPUs. CPFloat supports a comprehensive set of rounding modes that makes it highly versatile for observing the numerical behavior of low-precision arithmetic. However, CPFloat is not designed for neural network quantization but rather for general-purpose low-precision arithmetic in numerical algorithms.
@@
QPyTorch (Zhang et al., 2019) is a PyTorch-based low-precision simulator, specifically designed for low-precision arithmetic for neural network training without relying on low-precision hardware. As mentioned in (Fasi and Mikaitis, 2023), its principles are analogous to those of chop, in that numbers are stored in binary32 before as well as after rounding, and offers RNE, RNZ, and SR rounding modes, enabling efficient training (e.g., CIFAR10) via fused CUDA kernels. However, infinities, NaNs, and subnormals are excluded for efficiency because neural network training typically doesn’t involve these special values. Besides, its limited rounding support (lacking RZ, RUD, RO) restricts quantization studies and numerical simulations.

### v5

@@ Abstract.
Motivated by the growing demand for reduced-precision arithmetic in computational science, we exploit lower-precision emulation in Python—widely regarded as the dominant programming language for numerical analysis and machine learning. Low-precision paradigms have revolutionized deep learning by enabling more efficient computation and reduced memory footprint while maintaining model fidelity. To better enable numerical experimentation with and exploration of reduced-precision computation, we developed the Pychop, which supports customizable floating-point formats and a comprehensive set of rounding modes in Python, allowing users to benefit from fast, reduced-precision emulation in numerous applications. Pychop also introduces interfaces for both PyTorch and JAX, enabling efficient reduced-precision emulation on GPUs for neural network training and inference with unparalleled flexibility.
In this paper, we offer a comprehensive exposition of the design and applications of Pychop, establishing it as a foundational tool for advancing mixed-precision algorithms. Furthermore, we present empirical results on reduced-precision emulation for image classification and object detection using published datasets, illustrating the sensitivity of the use of low precision and offering valuable insights into its quantization-aware training and post-quantization impacts. Pychop enables in-depth investigations into the effects of numerical precision, facilitates the development of novel hardware accelerators, and integrates seamlessly into existing deep learning workflows.
@@
$c$s: Mathematics of computing Mathematical software performanceccs: Computing methodologies Modeling and simulationccs: Software and its engineering
@@ 1. Introduction
The increasing support of reduced-precision arithmetic in hardware architectures has triggered a resurgence of interest in mixed-precision algorithms, particularly within the fields of numerical analysis and deep learning. There are numerous benefits to using numerical formats with lower precision than single precision primarily because they require less memory bandwidth. The mixed-precision algorithms, which strategically combine reduced-precision and high-precision computations, have emerged as a heating topic of research due to their capability to enhance algorithmic performance across various domains—most notably energy efficiency, data transfer, and arithmetic speedup—while maintaining acceptable levels of numerical accuracy and stability (Higham and Mary, 2022). This paradigm exploits the inherent trade-offs between computational cost and precision, offering a compelling technique for addressing the escalating demands of large-scale numerical simulations and machine learning applications. The ability to tailor precision to specific computational tasks not only reduces energy consumption but also reduces processing times, making mixed-precision computations indispensable in the era of exascale computing and resource-constrained environments.
It is known that the increasing size of neural networks typically improves model generalization ability and accuracy at the cost of memory and compute resources for model deployment. The advance of mixed-precision training has been a milestone for deep learning efficiency. Micikevicius et al. (2018) established its use in practice, demonstrating that fp16 (half-precision, 16 bits total with 5 exponent and 10 significand bits) could handle weights and activations while fp32 accumulations preserved gradient fidelity, yielding speedups of $2-3\times$ on NVIDIA Volta GPUs. This approach has been integrated into PyTorch (Paszke et al., 2019) and implemented via Automatic Mixed Precision (AMP) [^1] https://pytorch.org/docs/stable/amp.html, which automates training in low precision with gradient scaling, lowering memory consumption by approximately 50% for models like ResNet-50 (He et al., 2016). However, AMP’s dependence on hardware-supported fp16 hinders its flexibility to other floating-point formats, a constraint our emulator eliminates by supporting arbitrary precision configurations. Further, (Perez et al., 2023) propose the training method via the scalings for fp8 (8-bit floating-point with two formats (Noune et al., 2022; Micikevicius et al., 2022)—exponent bits and 2 significand bits for E5 and 4 exponent bits and 3 significand bits for E4) linear layers by dynamically updating per-tensor scales for the weights, gradients and activations, validating an acceptable performance of GPT and Llama 2 deployed with fp8 precision across model sizes ranging from  [^1] M to 70B.
@@
Table 1. Key parameters of seven floating-point formats; $u$ denotes the unit roundoff corresponding to the precision, $x_{\min}$ denotes the smallest positive normalized floating-point number, $x_{\max}$ denotes the largest floating-point number, $t$ denotes the number of binary digits in the significand (including the implicit leading bit), $e_{\min}$ denotes the exponent of $x_{\min}$, and $e_{\max}$ denotes the exponent of $x_{\max}$. The last two columns give the number of exponent bits and significand bits (excluding the implicit bit).
$u$ $x_{\min}$ $x_{\max}$ $t$ $e_{\min}$ $e_{\max}$ exp. bits sig. bits
NVIDIA quarter precision (e4m3) $6.25\times 10^{-2}$ $1.5625\times 10^{-2}$ $2.4\times 10^{2}$ 4 -6 7 4 3
NVIDIA quarter precision (e5m2) $1.25\times 10^{-1}$ $6.10\times 10^{-5}$ $5.73\times 10^{4}$ 3 -14 15 5 2
@@
TensorFloat-32 (tf32) $9.77\times 10^{-4}$ $1.18\times 10^{-38}$ $3.40\times 10^{38}$ 11 -126 127 8 10
@@
We design a reduced-precision emulation software called Pychop for Python, offering a highly customizable setting for users to simulate arbitrary reduced-precision arithmetic—floating-point, fixed-point, and integer—with a diverse of rounding modes. This emulator transcends the constraints of fixed hardware by allowing users to define custom precision formats—specifying exponent and significand bits—and to select from rounding modes, e.g., round to nearest, round up / down, round toward zero, round toward odd, and two stochastic variants (proportional and uniform probability). The software promotes comprehensive precision format support, flexibility, and a beginner-friendly interface:
  • Unmatched Flexibility: Emulating both standard (see Table 1) and customized numerical formats for reduced-precision emulation, enabling researchers to prototype hypothetical hardware or explore theoretical precision boundaries without physical constraints.
  • Precision Granularity: Providing precise control over numerical representation and supporting soft error emulation in arithmetic operation, which is critical for dissecting the impact of quantization on both numerical algorithms and deep learning methods, e.g., gradient updates, weight distributions, and activation ranges in neural networks.
  • Seamless Integration: Offering direct emulators into PyTorch layers, enabling practitioners to experiment with mixed-precision training and inference pipelines with minimal overhead. The user-friendly API enables users to deploy quantization-aware and post-quantization strategies easily.
  • Rounding Mode Exploration: Supporting a diverse collection of rounding modes is available, enabling empirical evaluation of their roles in numerical stability. Our stochastic rounding is high-performance implementation, which shows empirically comparable performance to deterministic rounding.
Using Pychop, we present a comprehensive evaluation of training neural networks with quantization-aware training and post-quantization strategies in image classification tasks and object detection tasks, which offer insights into how different precision with rounding modes behave with respect to performance gains. Our software and experimental code are publicly available at https://github.com/inEXASCALE/pychop.
@@ 2. Related Work
Table 2. Software for simulating reduced-precision arithmetic
@@
FaultChop (Higham and Pranesh, 2019) MATLAB fp32/fp64 R S F fp16/bf16 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
gfloat (Research, 2024) Python custom A A F fp16/bf16/e5m2/e4m3/OCP/MX $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Pychop Python / MATLAB fp32/fp64 R S F e5m2/e4m3/fp16/bf16/tf32 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Low-precision arithmetic has emerged as a pivotal technique for optimizing computational efficiency in scientific computing and machine learning applications, where reduced precision can significantly lower resource demands while maintaining acceptable accuracy (Higham and Pranesh, 2019). Several software libraries have been developed to emulate reduced-precision arithmetic, each with distinct capabilities tailored to specific use cases. In this section, we discuss the software packages listed in Table 2222The table’s design follows (Fasi and Mikaitis, 2023), highlighting their strengths and limitations in the context of reduced-precision arithmetic simulation.
@@
GNU MPFR (Fousse et al., 2007) is a C library for simulating multiple-precision floating-point computations with guaranteed correct rounding. It excels in scenarios requiring arbitrary precision for both the significand ($p$) and exponent range ($e$). It is a preferred choice for numerical simulations demanding high accuracy, such as symbolic computation and numerical analysis. However, its default lack of subnormal number support ($s$, denoted as O) requires explicit user configuration, which can complicate workflows. Besides, GNU MPFR does not offer built-in floating-point formats and is not suited for neural network training, limiting its capability in reduced-precision emulation in deep learning.
@@
SIPE (Lefèvre, 2013), another C-based library, focuses on very reduced-precision computations with correct rounding. It supports multiple storage formats, restricted significand ($p$), and a restricted exponent range ($e$), while fully supporting subnormal numbers ($s$). SIPE implements RNE and RZ rounding modes for reduced-precision simulations which is suitable for exploring numerical stability and precision trade-offs in low-bitwidth computations, such as those encountered in embedded systems. Its primary limitation lies in its restricted rounding mode support and lack of built-in formats, which reduces its versatility.
@@
rpe (Dawson and Düben, 2017), implemented in Fortran, is tailored for emulating reduced floating-point precision in large-scale numerical simulations, such as climate modeling. It operates with restricted significand ($p$) and exponent range ($e$) limited to built-in types (B) and supports subnormal numbers only for built-in types (B). A key advantage is its native support for the fp16 format, aligning with reduced-precision hardware standards like IEEE 754 binary16. However, rpe’s exclusive support for RNE rounding limits its flexibility in scenarios requiring diverse rounding strategies, and it does not support neural network training, focusing solely on numerical simulations.
@@
FloatX (Flegar et al., 2019) and FlexFloat (Tagliavini et al., 2020) are C++ libraries that provide frameworks for customized floating-point arithmetic in reduced-precision simulations. These libraries support restricted significand ((p)) and exponent range ((e)), fully accommodate subnormal numbers ((s)), and utilize fp32/fp64 storage formats for compatibility with standard representations. Their simplicity, limited to round-to-nearest-even (RNE) rounding, makes them accessible for educational purposes and prototyping. However, this restricted rounding support limits their adaptability, and neither library includes predefined formats or supports neural network training, confining their use to general-purpose reduced-precision arithmetic experimentation. Both FloatX and FlexFloat strictly follow standard C++ conventions, including round-to-nearest, ties-to-even rounding, and default datatype casting. Although these conventions ensure compatibility with native floating-point operations, they limit users’ ability to explore diverse rounding strategies in numerical simulations. This constraint poses a notable challenge for researchers and practitioners needing flexible, application-specific rounding behaviors.
@@
INTLAB (Rump, 1999), a MATLAB toolbox, leverages interval arithmetic to facilitate reduced-precision floating-point simulation. It uses fp64 storage, supports restricted significand ($p$) and exponent range ($e$), and fully supports subnormal numbers ($s$). INTLAB provides RNE, RZ, and RUD rounding modes, offering moderate flexibility for numerical computations in MATLAB environments, such as verified computing. Its lack of built-in formats and optimization for neural network training limits its scope, positioning it as a tool for reliable numerical analysis rather than machine learning.
@@
FaultChop (Higham and Pranesh, 2019), another MATLAB library, enables reduced-precision arithmetic simulation with fp32/fp64 storage, restricted significand ($p$), and exponent range ($e$), alongside flexible subnormal number support (F). It supports built-in fp16 and bf16 formats, aligning with modern hardware standards, and implements RNE, RZ, RUD, and stochastic rounding (SR). The inclusion of SR is particularly valuable for simulating quantization effects, which are critical in machine learning research. Despite this, chop does not directly support neural network training and inference, and it is limited to the MATLAB environment, restricting its application to numerical experimentation and analysis of quantization impacts in machine learning.
@@
CPFloat (Fasi and Mikaitis, 2023) is a C library optimized for efficient reduced-precision arithmetic simulation, supporting fp32/fp64 storage with restricted significand ($p$), and exponent range ($e$). It includes built-in fp16, bf16, and tf32 formats with flexible subnormal number support (F), enhancing compatibility with hardware-accelerated systems such as GPUs. CPFloat supports a comprehensive set of rounding modes that makes it highly versatile for observing the numerical behavior of reduced-precision arithmetic. However, CPFloat is not designed for neural network quantization but rather for general-purpose reduced-precision arithmetic in numerical algorithms.
@@
For reduced-precision emulation in Python, QPyTorch (Zhang et al., 2019) is a PyTorch-based reduced-precision simulator, specifically designed for reduced-precision arithmetic for neural network training without relying on reduced-precision hardware. As mentioned in (Fasi and Mikaitis, 2023), its principles are analogous to those of FaultChop, in that numbers are stored in binary32 before as well as after rounding, and offers RNE, RNZ, and SR rounding modes, enabling efficient training via fused CUDA kernels. However, infinities, NaNs, and subnormals are excluded for efficiency because neural network training typically doesn’t involve these special values. Further, its limited rounding mode support (lacking RZ, RUD, RO) restricts quantization studies and numerical simulations. gfloat (Research, 2024) is a Python library designed for simulating reduced-precision floating-point arithmetic. It allows experimentation with various floating-point formats in Python, including IEEE 754 and OCP/MX (Microscaling Formats) (Goldberg, 1991; Darvish Rouhani et al., 2023), and supports stochastic rounding, but does not include soft-error or fault-injection simulation features. gfloat supports array in NumPy, PyTorch, and JAX.